OpenStudy (anonymous):

Solve for r in the equation C = 2d – 2r

OpenStudy (anonymous):

c-2d=-2r r=(2d-c)/2

OpenStudy (anonymous):

C = 2d – 2r r=(2d-C)/2

OpenStudy (curry):

c-2d=-2r r=(2d-c)/2